OpenStudy (anonymous):

Consider the set D of all numbers in the interval (-infinity , 5). What is the complement of D ?

OpenStudy (anonymous):

I think it's (5, +infinity) . I've never see it .

OpenStudy (anonymous):

The first is the set: \[\{x\in R\phantom{.}|\phantom{.}x\in (-\infty,5)\}\] The complement would be the rest of the number line from \(5\) to \(+\infty\) which is notated as : \[\{x\in R\phantom{.}|\phantom{.}x\in (5,\infty)\}\]

OpenStudy (anonymous):

The choices are A.(-∞,-5) B.[5,+∞) C.(-∞,5) D.(5+∞)

OpenStudy (anonymous):

OOOOOH Well we know that \(5\) and \(+\infty\) are part of the interval so we are given the options of: \([5,+\infty)\) and \((5,+\infty)\) EXACTLY, how as the original set notated?

OpenStudy (anonymous):

I typed B & D the way it was on my paper . :c

OpenStudy (anonymous):

Haha no like the original set not the complement, the one from negative infinity to 5

OpenStudy (anonymous):

Like did it have two brackets like this: \((\phantom{...})\)?

OpenStudy (anonymous):

Everything that's on here is what's on the paper , & Yeah it was (-∞,5)

OpenStudy (mathstudent55):

Set D does not include5, so the complement must include 5: D': \([5, \infty) \) The answer is B.

OpenStudy (mathstudent55):